# A distance function for stochastic matrices

### v1

@@ II Sequence distances
Here $\lambda_{i}$ are the eigenvalues of $\boldsymbol{P}$, the $\boldsymbol{e}^{(i)}$ are an orthonormal basis $\boldsymbol{e}^{(i)}\cdot\boldsymbol{e}^{(j)}=\delta_{ij}$, with $\boldsymbol{e}^{(1)}=\boldsymbol{\pi}_{\star}^{\circ\frac{1}{2}}$. Finally the assumption $\boldsymbol{\pi}_{\star}>\boldsymbol{0}$ for the unique stationary distribution of $\boldsymbol{P}$ is imposed. Usefully, the eigenvalues of ergodic and reversible Markov chains are real and satisfy $\lambda_{1}=1>\lambda_{2}\geq\ldots\geq\lambda_{n}>-1$. This allows us to take the limit $\tau\rightarrow\infty$ and know with certainty the power of all eigenvalues, except $\lambda_{1}$, will eventually vanish. With the expression in Eq. 15, and using the Bhattacharyya angle’s Taylor series, with associated bounds on remainder terms, we can give two conditions needed to be sufficient to bound the mixing time,
@@
Where we have introduced the so-called spectral gap $\lambda_{\max}=\max(\lambda_{2},|\lambda_{n}|)$, with $\pi_{-}=\min\boldsymbol{\pi}_{\star}$, and the constant $\gamma=1-(5/16)^{2/7}\approx 0.282$. The constant $\gamma$ is chosen to ensure the remainder of the Taylor series expansion of the mixing time can be replaced safely with a known lower bound, which entails a minimum requirement on $\tau_{\star}$ given by $\tau_{-}$. The second inequality using $\tau_{+}$ gives a sufficient condition which guarantees the desired distance accuracy $\epsilon$ has been achieved. Consequently, the true mixing time is known to fall within the range $[\tau_{-},\tau_{+}]$. Given the number of uses, and easily obtained results, afforded by the Bhattacharyya angle we promote it as a robust tool to investigate both short and long time Markov chain runs.
@@ III Stochastic matrix distances
Our idea is to, again appealing to information geometry, elevate the Bhattacharyya angle from a probability space to matrices. To this end, note that any stochastic matrix is a collection of individual state-conditional next-state probability distributions [52, 53]. This can be interpreted as a stochastic matrix being an element of a Cartesian product space of underlying categorical distributions (as explained in the trace distance section). Thus we identify $\boldsymbol{P}=(\boldsymbol{p}_{1},\ldots,\boldsymbol{p}_{n})\in\Delta\times\cdots\times\Delta$, where $\Delta$ is the $n$ dimensional simplex space, i.e. $\boldsymbol{P}$ is an $n\times n$ matrix. Such a decomposition induces what is known as a product metric in differential geometry [41]. In other words, we can build a distance function on the space of stochastic matrices by using the Bhattacharyya angle inherited from the underlying simplex space $\Delta$. Using a product metric structure here is actually very natural. The work of [52, 53] characterise the metric tensor components for stochastic matrices in terms of invariance under linear transformations called Markov maps. The result is choosing a product metric structure or asking for invariance under Markov maps results in the same distance on stochastic matrices.
@@
$g^{\mathrm{prod}}_{ij}\left(\boldsymbol{p}_{1},\boldsymbol{p}_{2}\right)=g_{ij}\left(\boldsymbol{p}_{1}\right)\oplus g_{ij}\left(\boldsymbol{p}_{2}\right)$ (17)
@@
Further, we can construct the product curve which connects pairs of points on $\Delta\times\Delta$ as $\boldsymbol{c}(t)=\left(\boldsymbol{c}_{1}(t),\boldsymbol{c}_{2}(t)\right)$ where $\boldsymbol{c}_{1}(t)$ and $\boldsymbol{c}_{2}(t)$ are the geodesic curves on the two manifolds respectively. A well known result in differential geometry is that product metrics induce an associated distance measure on the product manifold,
$d_{\mathrm{prod}}\left(\boldsymbol{p}_{1}\oplus\boldsymbol{q}_{1},\boldsymbol{p}_{2}\oplus\boldsymbol{q}_{2}\right)=$ (18)
$\quad\sqrt{d^{2}_{1}\left(\boldsymbol{p}_{1},\boldsymbol{p}_{2}\right)+d^{2}_{2}\left(\boldsymbol{q}_{1},\boldsymbol{q}_{2}\right)}$
@@
More generally, for an $n$ dimensional simplex we have the general form for a distance measure as $d_{\mathrm{prod}}=\sqrt{\sum_{i=1}^{n}d^{2}_{i}}$ and, consequently, the distance function induced by using the product metric of the usual Bhattacharyya angle is,
$d\left(\boldsymbol{P}_{1},\boldsymbol{P}_{2}\right)=$ (19)
$\quad 2\sqrt{\sum_{x_{i}\in X}\arccos^{2}\sum_{x_{j}\in X}\sqrt{P_{1}(x_{i},x_{j})P_{2}(x_{i},x_{j})}}$
@@
As well as enjoying many of the previously defined properties of the Bhattacharyya angle, it is also valid for any pair of stochastic matrices. There are no other requirements to compute the distance on the space of stochastic matrices. This appears to be a genuinely new distance function derived via information geometry.
Although motivated by its use for short time Markov chains, we can again consider infinite step distances for certain classes of models. In the case of ergodic (irreducible and aperiodic) Markov chains, with respective stationary distributions $\boldsymbol{\pi}_{1}$ and $\boldsymbol{\pi}_{2}$, the stochastic matrix distance has the limit,
$\lim_{N\rightarrow\infty}d\left(\boldsymbol{P}_{1}^{N},\boldsymbol{P}_{2}^{N}\right)=2\sqrt{n}\arccos\mathrm{BC}(\boldsymbol{\pi}_{1},\boldsymbol{\pi}_{2})$ (20)
@@
Notice that $\lim_{N\rightarrow\infty}^{n\rightarrow\infty}d\left(\boldsymbol{P}_{1}^{N},\boldsymbol{P}_{2}^{N}\right)/\sqrt{n}$ is finite and equal to the Bhattacharyya angle between the stationary distributions of the Markov chains. This motivates the definition of a general distance function for ergodic Markov chains as,
$d_{E}(M_{1},M_{2})=2\arccos\mathrm{BC}(\boldsymbol{\pi}_{1},\boldsymbol{\pi}_{2})$ (21)
@@
Intuitively this makes sense as the long term run of an ergodic Markov chain converges to its stationary distribution. Therefore any differences between two Markov chains should only be accounted for by their stationary distributions, and not initial conditions. As hinted at from the mixing times section, the stochastic matrix distance for ergodic chains is also identical to considering the distance between the asymptotic distributions from starting in any arbitrary initial distribution, i.e. $d^{(n)}(\boldsymbol{e}_{x}^{T}\boldsymbol{P}^{N}_{1},\boldsymbol{e}_{y}^{T}\boldsymbol{P}^{N}_{2})\rightarrow d_{E}(M_{1},M_{2})$ as $N\rightarrow\infty$. This adds further weight to using the Bhattacharyya angle and stochastic matrix distance as they have a natural correspondence on ergodic chains.
@@ IV Numerical Simulation
Figure 1: Ternary plot of the four base Dirichlet distributions used to generate clusters of stochastic matrices. The top-left density plot serves as a reference distribution. The parameters $\boldsymbol{\alpha}_{i}(t)$ approach $\boldsymbol{\alpha}_{0}$ linearly such that all distributions eventually coincide, and thus cannot be distinguished.
@@
To illustrate the use of the new stochastic matrix distance, and compare it alongside other typical distance functions, we simulate the random generation of four clusters of stochastic matrices. For each cluster, we generate a set of $3\times 3$ stochastic matrices where each row is drawn from a Dirichlet distribution with parameter vector $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$. The vector $\boldsymbol{\alpha}$ is a strictly positive and real vector, and controls the centre and spread of the corresponding Dirichlet distribution. This setup means we can tune the centres and spread of each cluster while making each stochastic matrix “similar” in both row values and overall matrix values. Initially we fix a distinguished cluster as a reference and denote this with Dirichlet parameters $\boldsymbol{\alpha}_{0}$. The remaining three clusters are generated with distinct parameters $\boldsymbol{\alpha}_{i}$. Finally, we increment along the transformation $\boldsymbol{\alpha}_{i}(t)=(1-t)\boldsymbol{\alpha}_{i}+t\boldsymbol{\alpha}_{0}$ for different values of $t\in[0,1]$, observing the accuracy of clustering for various distance functions. A pseudo-procedure for the simulation is laid out in IV.
@@
In our setting, we compared the symmetric Kullback-Leibler divergence rate, matrix distances induced by the $L^{1}$ and $L^{2}$ norms, and our new stochastic matrix distances. We use cluster Dirichlet parameters, visualised in Fig. 1, $\boldsymbol{\alpha}_{0}=(8,8,8)$, $\boldsymbol{\alpha}_{1}=(8,2,2)$, $\boldsymbol{\alpha}_{2}=(2,8,2)$, $\boldsymbol{\alpha}_{3}=(2,2,8)$.
The results of the simulation are shown in Fig. 2. We observe that the metrics all have a similar behaviour. Moreover, the stochastic matrix distance performance is comparable to the other functions. All functions approach a mean adjusted Rand score of zero as the distributions for clusters converge. Notably the symmetric Kullback-Leibler divergence rate seems to perform the worst on average, but we do not emphasise this result particularly as we simulated well clustered data in a rigorous setting.
@@
[htbp] (): Steps to generate simulated data for testing different stochastic matrix distance functions. The Rand index is a measure of how well data has been clustered according to the true cluster label. Averaging the Rand index over many runs mitigates observing a single extreme result. Data: $\{\boldsymbol{\alpha}_{k}\}_{k=0}^{K}$ (cluster parameters), $T$ (time steps), $R$ (repetitions), $N$ (cluster elements), $m$ (distance function) Result: For function $m$, $K+1\geq 2$ clusters with $N$ elements each, output an array of pairs $(t,\bar{s})$ for step $t$, mean Rand index $\bar{s}$ over $R$ clustering repetitions $t\leftarrow 0$;
while $t\leq T$ do
       $\boldsymbol{\alpha}_{k}(t)\leftarrow(1-t)\boldsymbol{\alpha}_{k}+t\boldsymbol{\alpha}_{0}$;
       $r\leftarrow 0$;
       while $r\leq R$ do
             generate $K\times N$ stochastic matrices;
             with function $m$ calculate distance matrix for all pairs of matrices;
             cluster matrices into $K$ labels;
             calculate Rand index $s(r)$ for clustering accuracy;
             $r\leftarrow r+1$
       end while
      Calculate $\bar{s}(t)=\frac{1}{R}\sum_{r=1}^{R}s(r)$;
       $t\leftarrow t+\frac{1}{T}$
end while

### v2

@@ II Sequence distances
Here $\lambda_{i}$ are the eigenvalues of $\bm{P}$, the $\bm{e}^{(i)}$ are an orthonormal basis $\bm{e}^{(i)}\cdot\bm{e}^{(j)}=\delta_{ij}$, with $\bm{e}^{(1)}=\bm{\pi}_{\star}^{\circ\frac{1}{2}}$. Finally the assumption $\bm{\pi}_{\star}>\bm{0}$ for the unique stationary distribution of $\bm{P}$ is imposed. Usefully, the eigenvalues of ergodic and reversible Markov chains are real and satisfy $\lambda_{1}=1>\lambda_{2}\geq\ldots\geq\lambda_{n}>-1$. This allows us to take the limit $\tau\rightarrow\infty$ and know with certainty the power of all eigenvalues, except $\lambda_{1}$, will eventually vanish. With the expression in Eq. 16, and using the Bhattacharyya angle’s Taylor series, with associated bounds on remainder terms, we can give two conditions needed to be sufficient to bound the mixing time,
@@
where we have introduced the so-called spectral gap $\lambda_{\max}=\max(\lambda_{2},|\lambda_{n}|)$, with $\pi_{-}=\min\bm{\pi}_{\star}$, and the constant $\gamma=1-(5/16)^{2/7}\approx 0.282$. The constant $\gamma$ is chosen to ensure the remainder of the Taylor series expansion of the mixing time can be replaced safely with a known lower bound, which entails a minimum requirement on $\tau_{\star}$ given by $\tau_{-}$. The second inequality using $\tau_{+}$ gives a sufficient condition which guarantees the desired distance accuracy $\epsilon$ has been achieved. Consequently, the true mixing time is known to fall within the range $[\tau_{-},\tau_{+}]$. Given the number of uses, and easily obtained results afforded by the Bhattacharyya angle we promote it as a robust tool to investigate both short and long time Markov chain runs.
@@ III Stochastic matrix distances
Our idea is to, again appealing to information geometry, elevate the Bhattacharyya angle from a probability space to matrices. To this end, note that any stochastic matrix is a collection of individual state-conditional next-state probability distributions [53, 54]. This can be interpreted as a stochastic matrix being an element of a Cartesian product space of underlying categorical distributions (as explained in the trace distance section). Thus we identify $\bm{P}=(\bm{p}_{1},\ldots,\bm{p}_{n})\in\Delta\times\cdots\times\Delta$, where $\Delta$ is the $n$ dimensional simplex space, i.e. $\bm{P}$ is an $n\times n$ matrix. Such a decomposition induces what is known as a product metric in differential geometry [42]. In other words, we can build a distance function on the space of stochastic matrices by using the Bhattacharyya angle inherited from the underlying simplex space $\Delta$. Using a product metric structure here is actually very natural. The work of [53, 54] characterise the metric tensor components for stochastic matrices in terms of invariance under linear transformations called Markov maps. The result is choosing a product metric structure or asking for invariance under Markov maps gives the same distance on stochastic matrices.
@@
$g^{\mathrm{prod}}_{ij}\left(\bm{p}_{1},\bm{p}_{2}\right)=g_{ij}\left(\bm{p}_{1}\right)\oplus g_{ij}\left(\bm{p}_{2}\right)$ (18)
@@
Further, we can construct the product curve which connects pairs of points on $\Delta\times\Delta$ as $\bm{c}(t)=\left(\bm{c}_{1}(t),\bm{c}_{2}(t)\right)$ where $\bm{c}_{1}(t)$ and $\bm{c}_{2}(t)$ are the geodesic curves on the two manifolds respectively. A well known result in differential geometry is that product metrics induce an associated distance measure on the product manifold,
$d_{\mathrm{prod}}\left(\bm{p}_{1}\oplus\bm{q}_{1},\bm{p}_{2}\oplus\bm{q}_{2}\right)=\sqrt{d^{2}_{1}\left(\bm{p}_{1},\bm{p}_{2}\right)+d^{2}_{2}\left(\bm{q}_{1},\bm{q}_{2}\right)}$ (19)
@@
More generally, for an $n$ dimensional simplex we have the general form for a distance measure as $d_{\mathrm{prod}}=\sqrt{\sum_{i=1}^{n}d^{2}_{i}}$ and, consequently, we can define a distance function on the space of stochastic matrices.
Theorem 2 (Stochastic Matrix Distance).
Let $\bm{P}_{1}$, $\bm{P}_{2}$ be square (row) stochastic matrices. The minimum length geodesic distance between $\bm{P}_{1}$, $\bm{P}_{2}$ is
$d\left(\bm{P}_{1},\bm{P}_{2}\right)=2\sqrt{\sum_{x_{i}\in X}\arccos^{2}\sum_{x_{j}\in X}\sqrt{P_{1}(x_{i},x_{j})P_{2}(x_{i},x_{j})}}$ (20)
@@
The proof of theorem 2 can be found in appendix A. As well as enjoying many of the previously defined properties of the Bhattacharyya angle, it is also valid for any pair of stochastic matrices. There are no other requirements to compute the distance on the space of stochastic matrices. This appears to be a genuinely new distance function derived via information geometry. Although motivated by its use for short time Markov chains, we can again consider infinite step distances for ergodic Markov chains and can actually prove the distance reduces to the Bhattacharyya angle between the stationary distributions of the two chains,
Theorem 3 (Ergodic distance).
Let $M_{1}$, $M_{2}$ be two ergodic discrete time Markov chains with unique stationary distributions $\bm{\pi}_{1}$, $\bm{\pi}_{2}$ respectively. The ergodic distance between $M_{1}$ and $M_{2}$ is
$d_{E}(M_{1},M_{2})=2\arccos\mathrm{BC}(\bm{\pi}_{1},\bm{\pi}_{2})$ (21)
Proof.
In the case of a ergodic (i.e. irreducible and aperiodic) Markov chain $\bm{P}_{1}$, with stationary distribution $\bm{\pi}_{1}$, via the Perron-Frobenius theorem the matrix power of the model’s stochastic matrix has the limit,
$\lim_{N\rightarrow\infty}\bm{P}_{1}^{N}=\textbf{1}^{T}\bm{\pi}_{1}$ (22)
Such a limit is just a statement that the matrix power converges to a matrix where each row is a copy of the stationary distribution. From here it is straightforward to show for the two models
$\sum_{x_{j}\in X}\sqrt{P_{1}(x_{i},x_{j})P_{2}(x_{i},x_{j})}=\mathrm{BC}(\bm{\pi}_{1},\bm{\pi}_{2})$ (23)
As the above is independent of $x_{i}$, the first summation inEquation 20 reduces to
$|X|\arccos^{2}\mathrm{BC}(\bm{\pi}_{1},\bm{\pi}_{2})$ (24)
For Markov chains with finite states we have $|X|=n$ and thus obtain,
$\lim_{N\rightarrow\infty}d\left(\bm{P}_{1}^{N},\bm{P}_{2}^{N}\right)=2\sqrt{n}\arccos\mathrm{BC}(\bm{\pi}_{1},\bm{\pi}_{2})$ (25)
@@
Finally, notice that $\lim_{N\rightarrow\infty}^{n\rightarrow\infty}d\left(\bm{P}_{1}^{N},\bm{P}_{2}^{N}\right)/\sqrt{n}$ is finite, independent of the value of $n$ and thus equal to the Bhattacharyya angle between the stationary distributions of the Markov chains. ∎
@@
Intuitively this makes sense as the long term run of an ergodic Markov chain converges to its stationary distribution. Therefore any differences between two Markov chains should only be accounted for by their stationary distributions, and not initial conditions. As noted in the mixing times section, the stochastic matrix distance for ergodic chains is also identical to considering the distance between the asymptotic distributions from starting in any arbitrary initial distribution, i.e. $d^{(n)}(\bm{e}_{x}^{T}\bm{P}^{N}_{1},\bm{e}_{y}^{T}\bm{P}^{N}_{2})\rightarrow d_{E}(M_{1},M_{2})$ as $N\rightarrow\infty$. This adds further weight to using the Bhattacharyya angle and stochastic matrix distance as they have a natural correspondence on ergodic chains.
@@ IV Numerical Simulation
To illustrate the use of the new stochastic matrix distance, and compare it alongside other typical distance functions, we simulate the random generation of four clusters of stochastic matrices. For each cluster, we generate a set of $3\times 3$ stochastic matrices where each row is drawn from a Dirichlet distribution with parameter vector $\bm{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$. The vector $\bm{\alpha}$ is a strictly positive and real vector, and controls the centre and spread of the corresponding Dirichlet distribution. This setup means we can tune the centres and spread of each cluster while making each stochastic matrix “similar” in both row values and overall matrix values. Initially we fix a distinguished cluster as a reference and denote this with Dirichlet parameters $\bm{\alpha}_{0}$. The remaining three clusters are generated with distinct parameters $\bm{\alpha}_{i}$. Finally, we increment along the transformation $\bm{\alpha}_{i}(t)=(1-t)\bm{\alpha}_{i}+t\bm{\alpha}_{0}$ for different values of $t\in[0,1]$, observing the accuracy of clustering for various distance functions. A pseudo-procedure for the simulation is laid out in 1. We choose the (adjusted) Rand index as the figure of merit for the clustering accuracy and average the index over many runs to mitigate observing a single extreme result [55]. Intuitively, the Rand index measures agreement, across two given clusterings, by checking how many pairs of clustered data elements are in the same cluster, and which pairs of clustered data elements are not in the same cluster. Ideally, all pairings according to one clustering match all parings in the other cluster, giving complete agreement. In practice it is better to adjust the Rand index under the null hypothesis assumption the clusterings were randomly assigned with equal probability by permuting the data elements, hence giving the adjusted Rand index as a measure of agreement.
Formally, given $N$ data points $X=\{x_{i}|i\in[N]\}$, for two clusterings $U=\{U_{1},\ldots,U_{R}\}$ and $V=\{V_{1},\ldots,V_{C}\}$ we can quantify the two clusterings contingency as $n_{ij}=|U_{i}\cap V_{j}|$. The collection of integers $n_{ij}$ count the overlap between different clusters. In consequence, we define the auxiliary quantities,
$n_{i+}=\sum_{j=1}^{C}n_{ij},\quad n_{+j}=\sum_{i=1}^{R}n_{ij}$ (26a)
$\bar{t}=\sum_{i=1}^{R}\sum_{j=1}^{C}n^{2}_{ij},\quad\bar{r}=\sum_{i=1}^{R}n^{2}_{i+},\quad\bar{c}=\sum_{j=1}^{C}n^{2}_{+j}$ (26b)
$\begin{gathered}a=\frac{\bar{t}-N}{2},\quad b=\frac{\bar{r}-\bar{t}}{2},\quad c=\frac{\bar{c}-\bar{t}}{2},\\ d=\frac{\bar{t}-\bar{r}-\bar{c}+N^{2}}{2}\end{gathered}$ (26e)
Taken together, these auxiliary quantities represent the agreement between two different clusterings of the data $X$. We can collect the above into a matrix $\bm{M}$ and express the adjusted Rand index in a compact manner as,
$\bm{M}=\begin{pmatrix}a&b\\ c&d\end{pmatrix},$ $\bm{P}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ (27)
$\mathrm{Rand_{adjusted}}=\frac{\binom{N}{2}\mathrm{tr}\bm{M}-\mathrm{tr}\bm{P}\bm{M}^{2}}{\binom{N}{2}^{2}-\mathrm{tr}\bm{P}\bm{M}^{2}}$ (28)
@@
Where $\bm{P}$ is known as the $2\times 2$ exchange matrix (i.e. it exchanges rows and columns). We compared the symmetric Kullback-Leibler divergence rate, matrix distances induced by the $L^{1}$ and $L^{2}$ norms, and our new stochastic matrix distances. We use cluster Dirichlet parameters, visualised in Fig. 1, $\bm{\alpha}_{0}=(8,8,8)$, $\bm{\alpha}_{1}=(8,2,2)$, $\bm{\alpha}_{2}=(2,8,2)$, $\bm{\alpha}_{3}=(2,2,8)$.
The results of the simulation are shown in Fig. 2, where a value of one on the vertical axis indicates clusters have been assigned correctly, and lower values mean cluster memberships are incorrect. We observe that the metrics all have a similar behaviour. Moreover, the stochastic matrix distance performance is comparable to the other functions. All functions approach a mean adjusted Rand score of zero as the distributions for clusters converge i.e. the clusters are indistinguishable and hence cluster membership can not be determined. We stress that the purpose of this simulation is not to pick a “best” distance to perform clustering, as this is too subjective, but to show our new distance function indeed behaves sensibly in a controlled setting.
Figure 1: Ternary plot of the four base Dirichlet distributions used to generate clusters of stochastic matrices. The top-left density plot serves as a reference distribution. The parameters $\bm{\alpha}_{i}(t)$ approach $\bm{\alpha}_{0}$ linearly such that all distributions eventually coincide, and thus cannot be distinguished.
@@
1
Require : Cluster distribution parameters $\{\bm{\alpha}_{k}\}_{k=0}^{K}$ (which serve as the true labels for generated data), $T$ (time steps), $R$ (repetitions), $N$ (cluster size), $m$ (distance function)
Ensure : For function $m$, $K+1\geq 2$ clusters with $N$ members each, output an array of pairs $(t,\bar{s})$ for step $t$, mean Rand index $\bar{s}$ over $R$ clustering repetitions
2
3.
[…]
5 while $t\leq T$ do
6       Update cluster parameters for time t
7       $\alpha_{k}(t)\leftarrow(1-t)\alpha_{k}+t\alpha_{0}$
8       $r\leftarrow 0$
9       while $r<R$ do
10             Generate $K\cdot N$ stochastic matrices based on parameters $\{\alpha_{k}(t)\}_{k=0}^{K}$
11             Using function $m$, calculate the distance matrix for all pairs of the generated matrices
12             Cluster the $K\cdot N$ matrices into $K$ labels using the distance matrix
13             Calculate the Rand index $s(r)$ comparing the clustering labels to the known generative clusters
14             $r\leftarrow r+1$
15
16       end while
17      Calculate the mean (adjusted) Rand index for time $t$: $\bar{s}(t)\leftarrow\frac{1}{R}\sum_{r=0}^{R-1}s(r)$
18       $t\leftarrow t+1$
19
20 end while
Algorithm 1 Steps to generate simulated data for testing different stochastic matrix distance functions. The Rand index is a measure of how well data has been clustered according to the true cluster label (which is the cluster distribution parameter $\alpha_{k}$)